OpenStudy (anonymous):

Given the parent function of f(x) = x4, what change will occur when the function is changed to f, bracket one half x end bracket?

OpenStudy (anonymous):

A. Graph opens the same way and is narrower B. Graph opens the same way and is wider C. Graph opens the opposite way and is narrower D. Graph opens the opposite way and is wider

OpenStudy (princeharryyy):

Graph opens the same way and is wider

OpenStudy (princeharryyy):

f(1/2 *x) = (1/16)*x^4

OpenStudy (princeharryyy):

so, B

OpenStudy (anonymous):

thank u

OpenStudy (princeharryyy):

:)

OpenStudy (anonymous):

do you have time to help me with one more

OpenStudy (princeharryyy):

sure :)

OpenStudy (anonymous):

What is the end behavior of the function f(x) = 3x4-x3 + 2x2 + 4x + 5? Up on the left, up on the right Up on the left, down on the right Down on the left, up on the right Down on the left, down on the right

OpenStudy (princeharryyy):

Up on the left, up on the right

OpenStudy (anonymous):

thank you

OpenStudy (princeharryyy):

anything else

OpenStudy (anonymous):

if you don't mind, i don't want to impose on anything

OpenStudy (princeharryyy):

It's ok, I've few minutes.

OpenStudy (anonymous):

What are the possible numbers of positive, negative, and complex zeros of f(x) = 3x4 - 5x3 - x2 - 8x + 4? Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0 Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 Positive: 3 or 1; negative: 1; nomplex: 2 or 0 Positive: 4 or 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0

OpenStudy (princeharryyy):

Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0

OpenStudy (anonymous):

so A?

OpenStudy (princeharryyy):

number of sign changes are two, so maximum two positive roots are possible and others follow according to them. Yeah, so, A

OpenStudy (anonymous):

honestly i have several more to go but if you want i can ask someone else XD I don't want to drive you mad

OpenStudy (princeharryyy):

I have 5 minutes left.

OpenStudy (anonymous):

ok so maybe one more?

OpenStudy (princeharryyy):

you can hurry.

OpenStudy (princeharryyy):

ok

OpenStudy (anonymous):

What are the coordinates of the turning point for the function f(x) = (x + 3)3 + 1? (-3, -1) (-3, 1) (3, -1) (3, 1)

OpenStudy (princeharryyy):

(-3, 1)

OpenStudy (anonymous):

awesome thank you so much!!!!

OpenStudy (princeharryyy):

2 minutes left. you can ask 1 more if you want.

OpenStudy (anonymous):

Which of the following graphs represents the function f(x) = x3 - x2 - 10x - 8? graph with 3 real zeros, down on left, up on right graph with 3 real zeros, up on left, down on right graph with 4 real zeros, up on left, up on right graph with 4 real zeros, down on left, down on right

OpenStudy (princeharryyy):

3 real zeroes

OpenStudy (anonymous):

A or B?

OpenStudy (princeharryyy):

A

OpenStudy (princeharryyy):

I was checking that.!

OpenStudy (anonymous):

oh sorry XD

OpenStudy (princeharryyy):

So, Bieee , Won't be able to help with anymore.

OpenStudy (anonymous):

ok byyyeee thank you for everything

OpenStudy (princeharryyy):

u r welcome :)